Gabrielle just won $2.5 million in the state lottery. She is given the option of receiving a total of $1.3 million now, or she can elect to be paid $100,000 at the end of each of the next 25 years. If Gabrielle can earn 5% annually on her investments, from a strict economic point of view which option should she take and what formula to use?
Answer #1

PV of first option =1300000
PV of second =PMT*((1-(1+r)^-n)/r) =100000*((1-(1+5%)^-25)/5%) =1409394.46

Hence PV of second option is higher hence second option should be chosen